OpenStudy (anonymous):

what is the domain of g(x)= square root of x + 1?

OpenStudy (anonymous):

Do you mean \[g(x) = \sqrt{x+1}\] ? (I love how you can make all the fancy math symbols on this thing)

OpenStudy (anonymous):

yes, i just can't figure out how to make the math symbols on here.

hartnn (hartnn):

\(\ddot \smile\)

OpenStudy (anonymous):

Click the little "Equation" button

OpenStudy (anonymous):

\[g(x)=\sqrt{x+1}\] domain of a function means the range of x that can be taken. by looking at this function, we notice that RHS of the equal sign, there's a square root. now in order to make this square root meaningful, x+1 must satisfy the condition that x+1>= 0

OpenStudy (anonymous):

so it would be all real numbers greater than zero right?

OpenStudy (anonymous):

no. it must satisfy X + 1 >= 0. you have to solve this inequality for the range of x. x + 1 >= 0 x >= -1 so domain of this is any number that bigger or equal to -1

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

no worries